OpenStudy (anonymous):

sin3x = cos3x for 0<=x<=360

OpenStudy (yrelhan4):

You know the formula for sin3x and cos3x?

OpenStudy (anonymous):

nope

OpenStudy (yrelhan4):

3sin(x) – 4(sin(x))^3 4cos^3x - 3cosx.. equate this..

OpenStudy (anonymous):

i dont understand :(

OpenStudy (yrelhan4):

oh sorry. i meant sin3x=3sin(x) – 4(sin(x))^3 cos3x=4cos^3x - 3cosx..

OpenStudy (anonymous):

ok can you please solve it for me now?

OpenStudy (yrelhan4):

you just need to equate them and evaluate. Do it and tell me where you are getting stuck. I'll help you. 3sin(x) – 4(sin(x))^3 = 4cos^3x - 3cosx You just need to solve this..

OpenStudy (anonymous):

do we need to find the values of x?

OpenStudy (anonymous):

3sin(x) – 4(sin(x))^3 = 4cos^3x - 3cosx ok so i am up to here. What do i do next?

OpenStudy (yrelhan4):

reaarange them.. 3(sinx + cosx) = 4[sin^3(x) + cos^3(x)] .. to the RHS apply the formula for a^3 + b^3= (a+b)(a^2 - ab + b^2)..

OpenStudy (anonymous):

@yrelhan4 what do we need to find in this ques?

OpenStudy (yrelhan4):

value of x, maybe.. @niksva

OpenStudy (yrelhan4):

wait. i'll see.

OpenStudy (anonymous):

then buddy can i suggest u a different method if u guys dont mind

OpenStudy (anonymous):

yes please

OpenStudy (yrelhan4):

Go ahead.. and 3(sinx + cosx)=4[(six+cosx) (sin^2x - sinxcosx + cons^2x) 3/4(sinx + cosx)=[(six+cosx) (1-sinxcosx)] 3/4=(1-sinxcosx) sinxcosx= 1-(3/4) sinxcosx=1/4 sin2x=1/8 that should be it..

OpenStudy (anonymous):

can we write cos3x = sin(pi/2 - 3x) ?

OpenStudy (anonymous):

first step is sin3x = cos3x second step would be sin3x = sin (pi/2 - 3x) is that clear?

OpenStudy (anonymous):

@yrelhan4 can you now solve sin2x = 1/8?

OpenStudy (yrelhan4):

oh dang. this is way easier than i made it look. sorry about that.. divide by cos3x on the both sides.. tan3x=1 3x=arctan1 3x=45 x=15.. sigh.

OpenStudy (yrelhan4):

and @niksva method is right too. follow any of these. but ignore what i did earlier.

OpenStudy (anonymous):

first method is used in those problems where u have to simplify the expressions hope this will help

OpenStudy (anonymous):

thanks guys :)